OpenStudy (anonymous):

Harder Question But Its Easy ! Which of the following is the slope-intercept form of the equation 8x + y = -1?

OpenStudy (anonymous):

y=-17-8x -8 is the slope and -17 is the y- intercept

OpenStudy (anonymous):

these are my answer choices A. y = -8x − 1 B. y = -8x + 1 C. y = 8x − 1 D. y = 8x + 1

OpenStudy (anonymous):

i am sorry my eyes are going bad. I thought the -1 was -17 the answer is A

OpenStudy (anonymous):

lols okok